OpenStudy (anonymous):

f(x)=(x-1)(x^2+2)^3, find f'(x)

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

I used product rule but my answer isn't one of the multiple choice options :S

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

show your steps using the equation editor.

OpenStudy (anonymous):

you have to use product rule and chain rule

OpenStudy (anonymous):

yeah i did that and i got

OpenStudy (anonymous):

\[(x-1)(x^2+2)^3 \times 2x + (x^2+2)^3 \times 1\]

OpenStudy (anonymous):

ok, lets split this: \[ u=x-1\implies u'=1\\ v=(x^2+2)^3\implies v'=3(x^2+3)^2(2x)\\ f(x)=uv\\ f'(x)=u'v+uv' \]

OpenStudy (anonymous):

correction.. \[ u=x-1\implies u'=1\\ v=(x^2+2)^3\implies v'=3(x^2+\color{red}2)^2(2x)\\ f(x)=uv\\ f'(x)=u'v+uv' \]

OpenStudy (anonymous):

see what you missed there?

OpenStudy (anonymous):

your derivative of "v" was incorrect

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

yw